SHORT-WAVELENGTH (30 TO 150 μm) OPTICALLY PUMPED LASER LINES IN A METHANOL LASER

\underline{E. C. C. VASCONCELLOS}, Instituto de Física ”Gleb Wataghin,” Departamento de Eletrônica Quântica, Universidade Estadual de Campinas (UNICAMP), 13083-970 Campinas, SP, Brazil;

L. R. ZINK and K. M. EVENSON, Time and Frequency Division, National Institute of Standards and Technology, Boulder, CO 80303-3328.

A new short-wavelength far-infrared laser^{a} is used to generate optically pumped far-infrared laser radiation in the wavelength range 30 to 150 μm in methanol. The large-amplitude torsional motion in methanol accounts for its rich far-infrared spectra. Its infrared absorption bands overlap well with the CO_{2} laser providing for the generation of strong high-frequency laser lines in the far-infrared. Methanol is an excellent short wavelength laser medium and about 40% of its lines have wavelengths shorter than 150 μm. We have already obtained a four-fold increase in the number of laser lines in the 50 μm region, and are now pushing to increase the density of laser lines with λ < 50 μm. A list of the strong laser lines observed in methanol, in the range 30 to 150 μm, will be presented in this work.

aFor detailed description of the laser see paper presented by K.M. Evenson at this meeting.